if the mean of a random sample of size 400 is used to estimate the mean of an infinite population with a standard deviation of 60, the probability is 0.95 that the error is less than

5.88.

117.6.

1.96.

3.92.
Answer #1

Solution :

Given that,

Z/2 = 1.96

Margin of error = E = Z/2* ( /n)

= 1.96 * (60 / 400)

= 5.88

The probability is 0.95 that the error is less than 5.88